# Student member, IEEE

### v1

@@ 3.3 The multi-feature learning module
In this section, we introduce the multi-feature learning submodule from two aspects: The structure and the training method.
@@ 3.3.1 Structure
As shown in Fig.1, the module consists of five function networks and a context network. The function network_1 learns a nonlinear mapping between flight conditions $\vec{\mathbf{x}}_{1}$ and $C_{P}$, the function network_2 learns a nonlinear mapping between coordinate point $\vec{\mathbf{x}}_{2}$ on a wing surface and $C_{P}$, the function network_3 learns a nonlinear mapping between Riemannian metric $g_{ij}$ at coordinate point $\vec{\mathbf{x}}_{2}$ and $C_{P}$, the function network_4 learns a nonlinear mapping between connection $\mit\Gamma_{ij}^{k}$ at coordinate point $\vec{\mathbf{x}}_{2}$ and $C_{P}$, the function network_5 learns a nonlinear mapping between scalar curvature $S$ at coordinate point $\vec{\mathbf{x}}_{2}$ and $C_{P}$. The context network learns the fusion strategy (i.e., weight vector) for outputs generated by the five function networks.
The input of this module can be expressed as
$\begin{cases}\vec{\mathbf{x}}_{1}=[Ma,AoA,Re]\\ \vec{\mathbf{x}}_{2}=\left\{x,y,z\right\}\\ \vec{\mathbf{x}}_{3}=\left\{g_{ij}\right\}\\ \vec{\mathbf{x}}_{4}=\left\{\mit\Gamma_{ij}^{k}\right\}\\ \vec{\mathbf{x}}_{5}=\left\{S\right\}\end{cases}$ (12)
where $Ma$ denotes the incoming Mach number, $AoA$ denotes the Angle of Attack, $Re$ denotes the Reynolds number.
@@
$\hat{C}_{P}=\sum_{z=1}^{5}f_{z\alpha}(\vec{\mathbf{x}}_{z})\times c_{\left(z-1\right)K+\alpha}(\vec{\bm{\xi}})$ (13)
where $\hat{C}_{P}$ denotes the predicted value, $f_{z\alpha}(\vec{\mathbf{x}}_{z})$ denotes the $\alpha$th component of the output vector from function network_$z$, $K$ denotes the number of output nodes in function networks, $c_{\alpha}(\vec{\bm{\xi}})$ denotes the $\alpha$th component of the output vector from the context network, and $\vec{\bm{\xi}}=[\vec{\mathbf{x}}_{1},\vec{\mathbf{x}}_{2},\vec{\mathbf{x}}_{3},\vec{\mathbf{x}}_{4},\vec{\mathbf{x}}_{5}]$ denotes the concatenation of $\vec{\mathbf{x}}_{1}\sim\vec{\mathbf{x}}_{5}$.
@@ 3.3.2 Training method
The loss function of the multi-feature learning module is:
@@
where $M$ denotes the number of data points in training set, $C_{P_{tr}}^{\beta}$ denotes the real $C_{P}$ value of the $\beta$th sample in the training set, and $\hat{C}_{P}^{\beta}$ denotes the predicted value of the $\beta$th sample in the training set.
The training algorithm of the multi-feature learning module is shown as follows.
@@
1 Initialize the learning rate $\eta$, the parameter matrix of function network $\bm{\theta}_{f}$, the parameter matrix of context network $\bm{\theta}_{c}$;
2 for number of training iterations do
3       $\vec{\bm{\xi}}\leftarrow[\vec{\mathbf{x}}_{1},\vec{\mathbf{x}}_{2},\vec{\mathbf{x}}_{3},\vec{\mathbf{x}}_{4},\vec{\mathbf{x}}_{5}]$;
4       $\hat{C}_{P}\leftarrow\sum_{z=1}^{5}f_{z\alpha}(\vec{\mathbf{x}}_{z})\times c_{\left(z-1\right)K+\alpha}(\vec{\bm{\xi}})$;
5       $L\leftarrow\frac{1}{M}\sum_{\beta=1}^{M}\left(C_{P_{tr}}^{\beta}-\hat{C}_{P}^{\beta}\right)^{2}$;
6       $\bm{\theta}_{f}\leftarrow\bm{\theta}_{f}-\eta\frac{\partial L}{\partial\bm{\theta}_{f}}$;
7       $\bm{\theta}_{c}\leftarrow\bm{\theta}_{c}-\eta\frac{\partial L}{\partial\bm{\theta}_{c}}$;
8
9 end for
10return $\bm{\theta}_{f}$ and $\bm{\theta}_{c}$;
4 EXPERIMENTAL RESULTS AND ANALYSIS
In this section, five different models, namely Multi-task Learning (MTL) [48], Manifold-based Airfoil Geometric Feature Extraction and Discrepant Data Fusion Learning (MDF) [18], our method, MLP [49] and DAN [25] are compared in predicting $C_{P}$ on the wing surface of DLR-F11 aircraft [21].
@@ 4.2 Model settings
In this section, we introduce the structure and the input data of MTL, MDF, our method, MLP and DAN.
@@
Flight conditions, coordinates on the wing surface, and geometric features are different types of data, and their numerical value, distribution, and physical meaning are different. It has been demonstrated that using a context network to learn the weight vector to fuse $f_{z\alpha}(\vec{\mathbf{x}}_{z})$ outputs more accurate predicted coefficients than with directly concatenates $\vec{\mathbf{x}}_{1}\sim\vec{\mathbf{x}}_{5}$ as inputs [18, 19, 20]. In this paper, MTL, MDF and our method use a weight vector mechanism to predict $C_{P}$, MLP concatenates $\vec{\mathbf{x}}_{1}\sim\vec{\mathbf{x}}_{5}$ as inputs to predict $C_{P}$, and DAN concatenates features extracted from wing section images and flight conditions together as inputs to predict $C_{P}$.
Considering that the above models contain fully connected (FC) layers and convolutional (CV) layers, we design multiple structures for each model according to the number of FC layers, CV layers and hidden nodes. The optional numbers of CV and FC layers are 3 and 6, respectively. For FC layers, the optional numbers of hidden nodes are 16, 32, 64 and 128. For CV layers, the kernel size are $2\times 2$, the channel of each kernel is $2^{1+e}$, where $e$ denotes the sequence number of each CV layer ($e$ starts from 1). For MTL, MDF and our method, the number of hidden layers in each function network are the same (for example, if the function network_1 has 3 layers, the remaining function networks also have 3 hidden layers). Therefore, MTL, MDF and our method have 64 different structures, respectively. MLP and DAN have 8 different structures, respectively. Table 3 only introduce the structures with smallest test errors (the optimal structure).
@@
Our method additionally considers geometric features from 8 neighbors around point $P_{ab}$. Fig. 4 describes how to select the 8 neighbors on the DLR-F11 wing. Among the 9 points, the distance $d=0.01$. The input of function network_2 is a matrix (with shape $batchsize\times 1\times 9\times 3$) that contains 9 different $\vec{\mathbf{x}}_{2}$, therefore the function network_2 is a 2D CNN, which consists of three convolutional layers, each of which have a convolution kernel with shape $2\times 2\times 4$, $2\times 2\times 8$ and $2\times 2\times 16$, respectively. Similarly, the inputs of the function network_3 (with shape $batchsize\times 1\times 18\times 2$), function network_4 (with shape $batchsize\times 2\times 18\times 2$) and function network_5 (with shape $batchsize\times 9$) are also matrices contain 9 different $\vec{\mathbf{x}}_{3}$, $\vec{\mathbf{x}}_{4}$ and $\vec{\mathbf{x}}_{5}$, respectively. The structures of function network_1, function network_5 and context network are all set to “FC: $16\times 3$”.
The input data of MLP is exactly the same with our method. The MLP is a multi-layer perceptron with 3 layers and each layer has 128 nodes. The aim of comparing MLP with MTL, MDF and our method is to further verify the effectiveness of the weight vector mechanism. Because the source code of MLP is not published by [49], we implemented MLP for this experiment.
@@
activation function ReLU [51]
@@
Table 5: Average MSEs of MTL, MDF, our method, MLP and DAN. The average MSE reduction is calculated as the average value for each MSE reduction with different $AoA$.
$AoA$ MTL [48] MDF [18] Our method MLP [49] DAN [25]
MSE reduction
(our mehod V.S. DAN)
$1.49$ $1.53$ $\mathbf{1.46}$ $1.53$ $1.52$ 3.95%
12° $\mathbf{1.19}$ $1.21$ $1.23$ $1.21$ $1.21$ -1.65%
16° $8.26\times 10^{-1}$ $6.66\times 10^{-1}$ $6.54\times 10^{-1}$ $\mathbf{5.93\times 10^{-1}}$ $6.50\times 10^{-1}$0.62%
18° $3.31\times 10^{-1}$ $2.92\times 10^{-1}$ $\mathbf{2.03\times 10^{-1}}$ $2.22\times 10^{-1}$ $2.29\times 10^{-1}$ 11.35%
18.5° $2.22\times 10^{-1}$ $1.82\times 10^{-1}$ $1.50\times 10^{-1}$ $\mathbf{1.36\times 10^{-1}}$ $1.56\times 10^{-1}$ 3.85%
19° $1.71\times 10^{-1}$ $2.08\times 10^{-1}$ $\mathbf{7.72\times 10^{-2}}$ $1.99\times 10^{-1}$ $7.98\times 10^{-2}$ 3.26%
20° $1.70\times 10^{-1}$ $6.94\times 10^{-2}$ $\mathbf{1.41\times 10^{-2}}$ $1.82\times 10^{-2}$ $2.30\times 10^{-2}$ 38.70%
average $6.29\times 10^{-1}$ $5.94\times 10^{-1}$ $\mathbf{5.41\times 10^{-1}}$ $5.58\times 10^{-1}$ $5.53\times 10^{-1}$ 8.41%
@@
Figure 5: Training and validation loss corves for MTL, MDF, our method, MLP and DAN with $AoA=20$°.
@@ 4.3 Results and analysis
Figure 6: The predicted $C_{P}$ (a) and predicted errors (b) with $AoA=20$° and span width equals 15%. The horizontal axis denotes the x axis along the fuselage direction, and the vertical axis denotes the z axis along the fuselage normal direction.
Figure 7: The predicted $C_{P}$ (a) and predicted errors (b) with $AoA=20$° and span width equals 68.1%. The horizontal and vertical axes have the same meanings as depicted in Fig. 6.
@@
Table 5 displays the average test MSEs for MTL, MDF, our method, MLP and DAN. In every row, figures in bold represent the minimum test MSEs (all the test MSEs are not normalized). Generally, the test MSEs of $C_{P}$ predicted by all models decrease as the $AoA$ increases. This phenomenon occurs because the data samples in the large $AoA$ region are densely populated, whereas the data samples in the small $AoA$ region are relatively sparse. Comparing our method with DAN under the condition of $12$°$\leq AoA\leq 16$°, DAN exhibits smaller prediction MSEs than our method. However, for $AoA\geq 18$°, the prediction MSEs of DAN are bigger than those of our method. Similarly, when comparing our method with MLP, MLP shows smaller prediction MSEs at $AoA=16$° and $AoA=18.5$°, while our method exhibits smaller prediction MSEs for $AoA\geq 19$°. The above observations indicate that our method is more accurate in predicting $C_{P}$ than other models for large $AoA$. Conversely, MLP and DAN have advantages in the case of relatively small $AoA$.
Fig. 5 depicts the average training and validation corves for all the above models in the 7-fold experiments. The fluctuations in loss variations observed in subgraphs (a), (d) and (e) are larger than those in other subgraphs. This indicates a difference in the stability of the training processes. The models corresponding to subgraphs (b) and (c) (MDF and our method) are more stable. Furthermore, compared with subgraph (a), the models represented in subgraphs (b), (c), and (d) (MDF, our method, and MLP) converge faster than MTL. These three models all incorporate geometric features as part of their inputs, which suggests that the incorporation of geometric features accelerates the convergence of their training processes. It’s worth highlighting that the models corresponding to subgraphs (c) and (d) (our method and MLP) incorporate the geometric features from additional 8 neighbors as part of their inputs, which further accelerates the convergence of these two models.
@@
Figure 8: The predicted errors with $AoA=18$° (a) and $AoA=7$° (b), and span width equals 68.1%. The horizontal and vertical axes have the same meanings as depicted in Fig. 6.
Figure 9: The $C_{P}$ variations of test set, our method, MLP and DAN with $AoA=20$°. The horizontal axis denotes the x axis along the fuselage direction, and the vertical axis denotes $C_{P}$ value.
@@
Fig. 6 illustrates the predicted $C_{P}$ distributions and the corresponding prediction errors on the surface of DLR-F11 wing with $AoA=20$° and span width equals 15%. The DLR-F11 aircraft contains a three-stage wing, it is evident that the wing sections depicted in the graph are composed of three parts. In each row of every subgraph, the left part represents the first wing stage, the middle part represents the second wing stage, and the right part denotes the third wing stage. In subgraph (a), the $C_{P}$ distributions predicted by all models are similar, suggesting that each model is capable of learning the variations in $C_{P}$ distribution on the DLR-F11 wing surface. To clearly compare the details of the predicted $C_{P}$ values in subgraph (a), we calculate
@@
where $err^{\beta}$ denotes the prediction error of the $\beta$th sample, $C_{P_{te}}^{\beta}$ denotes the real $C_{P}$ value of the $\beta$th sample in the test set, and $\hat{C}_{P}^{\beta}$ denotes the predicted value of the $\beta$th sample in the test set. The $err^{\beta}$ of each model is plotted in subgraph (b). On the first wing stage, all the points generated by MTL and MDF show larger errors compared to the points generated by our method. At the leading edge on the first wing stage, the points generated by MTL, MDF, MLP and DNA also show larger errors than the points generated by our method. At the leading edge on the upper surface of the second wing stage, the points generated by MTL, MDF, MLP and DAN exhibit larger errors than the points generated by our method. The leading edge of a wing is the area where its $C_{P}$ changes dramatically, and our method is more accurate at the leading edge of both the first and second stages than other models. At the trailing edge on the upper surface of the second wing stage, the points generated by MTL, MDF, our method and DAN have larger errors than the points generated by MLP. On the lower surface of the second wing stage, the points generated by MTL, MDF, our method are not as accurate as the points generated by MLP and DAN. At the leading edge on the third wing stage, the points generated by MTL, MDF, MLP and DAN have larger errors than the points generated by our method. These observations demonstrate that our method is more accurate at the leading edge of wing sections than other methods.
@@
Fig. 7 illustrates the predicted $C_{P}$ distributions and the corresponding prediction errors on the surface of DLR-F11 wing with $AoA=20$° and span width equals 68.1%. Similar to Fig. 6, subgraph (a) demonstrates that all models can learn the variations of $C_{P}$ distributions. In subgraph (b), on the first wing stage, the points generated by MTL, MDF and MLP have larger errors than the points generated by our method and DAN. At the leading edge on the second wing stage, the points generated by MLP, DAN have larger errors than the points generated by MTL, MDF and our method. At the trailing edge of the second wing stage, one point generated by MDF is red. However, among all corresponding points generated by these comparative models, the points generated by our method have the darkest color, indicating that the $C_{P}$ error of this point predicted by our method is the smallest among the comparative models in this paper. On the third wing stage, the points generated by MTL show larger errors than the points generated by other methods. These observations demonstrate that our method is more accurate at the leading edge on the first wing stage than other methods.
@@
Figure 10: The $C_{P}$ variations of test set, our method, MLP and DAN with $AoA=18.5$°. The horizontal and vertical axes have the same meanings as depicted in Fig. 9.
@@
Fig. 8 displays the prediction errors of the DLR-F11 wing section with a span width of 68.1% at $AoA$ of 18° and 7°, respectively. In subgraph (a), at the leading edge on the first wing stage, the points generated by MDF, MLP and DAN show larger errors than the points generated by MTL and our method. The remaining stages are similar. In subgraph (b), it is difficult to see the differences among the models in detail. In addition, points on the upper surface of the second wing stage generated by all methods have large errors, because the DLR-F11 data samples in small $AoA$ region are sparser than those in large $AoA$. The observations in Fig. 6, Fig. 7 and Fig. 8 demonstrate that our method exhibits distinct advantages at the leading edge of wing sections in case with large $AoA$, while in cases with small $AoA$, our method closely aligns with existing methods, which is consistent with table 5.
@@
To better illustrate $C_{P}$ variations of different models, the $C_{P}$ variations of different wing sections at different $AoA$ are plotted in Fig. 9 and Fig. 10. It is worth noting that only three models (our method, MLP and DAN) are included to maintain clarity in graphs, as they exhibit relatively small prediction errors. In each subgraph, three set of curves are plotted to depict the $C_{P}$ variations on three stages of the DLR-F11 wing. The left curve represents the $C_{P}$ variation on the first wing stage, the middle curve represents the $C_{P}$ variation on the second wing stage, and the right curve represents the $C_{P}$ variation on the third wing stage. In each curve, the upper curve represents the $C_{P}$ variations on the lower surface of the DLR-F11 wing, and the lower curve represents the $C_{P}$ variations on the upper surface of the DLR-F11 wing.

### v2

@@ 3.2.1 Structure
The multi-feature learning module consists of five function networks and a context network.
The function network_1 learns a nonlinear mapping between flight conditions $\vec{\mathbf{x}}_{1}=[Ma,AoA,Re]$ and $C_{P}$, where $Ma$ denotes the incoming Mach number, $AoA$ denotes the Angle of Attack, $Re$ denotes the Reynolds number. The function network_2 learns a nonlinear mapping between coordinate $\vec{\mathbf{x}}_{2}=\left\{x,y,z\right\}$ on an aircraft and $C_{P}$. The function network_3 learns a nonlinear mapping between Riemannian metric $\vec{\mathbf{x}}_{3}=\left\{g_{ij}\right\}$ and $C_{P}$. The function network_4 learns a nonlinear mapping between connection $\vec{\mathbf{x}}_{4}=\left\{\mit\Gamma_{ij}^{k}\right\}$ and $C_{P}$. The function network_5 learns a nonlinear mapping between scalar curvature $\vec{\mathbf{x}}_{5}=\left\{S\right\}$ and $C_{P}$. The context network takes $\vec{\boldsymbol{\xi}}=[\vec{\mathbf{x}}_{1},\vec{\mathbf{x}}_{2},\vec{\mathbf{x}}_{3},\vec{\mathbf{x}}_{4},\vec{\mathbf{x}}_{5}]$ as input to learn weight vectors for function network outputs, where []delimited-[][$\cdot$][ ⋅ ] denotes the concatenation of vectors.
@@
$\hat{C}_{P}=\sum_{z=1}^{5}f_{z\alpha}(\vec{\mathbf{x}}_{z})\times c_{\left(z-1\right)K+\alpha}(\vec{\boldsymbol{\xi}})$ (9)
where $\hat{C}_{P}$ denotes the predicted value, $f_{z\alpha}(\vec{\mathbf{x}}_{z})$ denotes the $\alpha$th component of the output vector from function network_$z$, $K$ denotes the number of output nodes in each function networks, $c_{\alpha}(\vec{\boldsymbol{\xi}})$ denotes the $\alpha$th component of the weight vector generated by the context network.
@@ 3.2.2 Training Method
The loss function of this submodule is:
@@
where $M$ denotes the number of data points in training set, $C_{P_{tr}}^{\beta}$ denotes the real $C_{P}$ value of the $\beta$th sample in the training set, and $\hat{C}_{P}^{\beta}$ denotes the predicted value of the $\beta$th sample in the training set. The training algorithm of this module is shown as in Algorithm. 1.
@@
Parameter: learning rate $\eta$, function network parameters $\boldsymbol{\theta}_{f}$, context network parameters $\boldsymbol{\theta}_{c}$
Output: $\boldsymbol{\theta}_{f}$ and $\boldsymbol{\theta}_{c}$
1:  Initialize $\eta$, $\boldsymbol{\theta}_{f}$, $\boldsymbol{\theta}_{c}$, $e\leftarrow 0$
2:  while $e<$ training iterations do
3:     $\vec{\boldsymbol{\xi}}\leftarrow[\vec{\mathbf{x}}_{1},\vec{\mathbf{x}}_{2},\vec{\mathbf{x}}_{3},\vec{\mathbf{x}}_{4},\vec{\mathbf{x}}_{5}]$
4:     $\hat{C}_{P}\leftarrow\sum_{z=1}^{5}f_{z\alpha}(\vec{\mathbf{x}}_{z})\times c_{\left(z-1\right)K+\alpha}(\vec{\boldsymbol{\xi}})$
5:     $L\leftarrow\frac{1}{M}\sum_{\beta=1}^{M}\left(C_{P_{tr}}^{\beta}-\hat{C}_{P}^{\beta}\right)^{2}$
6:     $\boldsymbol{\theta}_{f}\leftarrow\boldsymbol{\theta}_{f}-\eta\frac{\partial L}{\partial\boldsymbol{\theta}_{f}}$
7:     $\boldsymbol{\theta}_{c}\leftarrow\boldsymbol{\theta}_{c}-\eta\frac{\partial L}{\partial\boldsymbol{\theta}_{c}}$
8:     $e\leftarrow e+1$
9:  end while
10:  return $\boldsymbol{\theta}_{f}$ and $\boldsymbol{\theta}_{c}$
4 Experiment and Analysis
In this section, five different models, namely Multi-task Learning (MTL) [35], Manifold-based Airfoil Geometric Feature Extraction and Discrepant Data Fusion Learning (MDF) [14], RGFil, MLP [36] and DAN [3] are compared in predicting $C_{P}$ on the wing of DLR-F11 aircraft [37]. Because the DLR-F11 dataset contains both geometry and flight condition parameters of an 3D aircraft, this dataset is chosen for testing all models compared in this paper.
@@ 4.2 Model settings
In this section, we introduce the structure and the input data of MTL, MDF, RGFiL, MLP and DAN.
@@
Flight conditions, coordinates on the wing surface, and geometric features are different types of data, and their numerical value, distribution, and physical meaning are different. It has been demonstrated that using a context network to learn the weight vector to fuse $f_{z\alpha}(\vec{\mathbf{x}}_{z})$ outputs more accurate predicted coefficients than with directly concatenates $\vec{\mathbf{x}}_{1}\sim\vec{\mathbf{x}}_{5}$ as inputs [14, 12, 13]. In this paper, MTL, MDF and RGFiL use a weight vector mechanism to predict $C_{P}$, MLP concatenates $\vec{\mathbf{x}}_{1}\sim\vec{\mathbf{x}}_{5}$ as inputs to predict $C_{P}$, and DAN concatenates features extracted from wing section images and flight conditions together as inputs to predict $C_{P}$.
Considering that the above models contain fully connected (FC) layers and convolutional (CV) layers, we design multiple structures for each model according to the number of FC layers, CV layers and hidden nodes. The optional numbers of CV and FC layers are 3 and 6, respectively. For FC layers, the optional numbers of hidden nodes are 16, 32, 64 and 128. For CV layers, the kernel size are $2\times 2$, the channel of each kernel is $2^{1+e}$, where $e$ denotes the sequence number of each CV layer ($e$ starts from 1). For MTL, MDF and RGFiL, the number of hidden layers in each function network are the same (for example, if the function network_1 has 3 layers, the remaining function networks also have 3 hidden layers). Therefore, MTL, MDF and RGFiL have 64 different structures, respectively. MLP and DAN have 8 different structures, respectively. Table 3 only introduce the structures with smallest test errors (the optimal structure).
@@
RGFiL additionally considers geometric features from 8 neighbors around point $P_{ab}$. Fig. 4 describes how to select the 8 neighbors on the DLR-F11 wing. The distance $d$ is a hyperparameter with optional values of 0.01, 0.005, and 0.001. The input of function network_2 is a matrix (with shape $batchsize\times 1\times 9\times 3$) that contains 9 different $\vec{\mathbf{x}}_{2}$, therefore the function network_2 is a 2D CNN, which consists of three convolutional layers, each of which have a convolution kernel with shape $2\times 2\times 4$, $2\times 2\times 8$ and $2\times 2\times 16$, respectively. Similarly, the inputs of the function network_3 (with shape $batchsize\times 1\times 18\times 2$), function network_4 (with shape $batchsize\times 2\times 18\times 2$) and function network_5 (with shape $batchsize\times 9$) are also matrices contain 9 different $\vec{\mathbf{x}}_{3}$, $\vec{\mathbf{x}}_{4}$ and $\vec{\mathbf{x}}_{5}$, respectively. The structures of function network_1, function network_5 and context network are all set to “FC: $16\times 3$”.
The input data of MLP is exactly the same with RGFiL. The MLP is a multi-layer perceptron with 6 layers and each layer has 128 nodes. The aim of comparing MLP with MTL, MDF and RGFiL is to further verify the effectiveness of the weight vector mechanism. Because the source code of MLP is not published by [36], we implemented MLP for this experiment.
@@
activation function LeakyReLU [39]
@@
Table 5: Average MSEs of MTL[35], MDF[14], RGFiL, MLP[36] and DAN[3]. $\eta$ denotes the MSE reduction of RGFiL_2 compared with DAN. The average MSE reduction is calculated as the average value for each MSE reduction.
$AoA$ MTL MDF
RGFiL_1
($d=0.01$)
RGFiL_2
($d=0.005$)
RGFiL_3
($d=0.001$)
MLP_1
($d=0.01$)
MLP_2
($d=0.005$)
MLP_3
($d=0.001$)
DAN $\eta$
1.49 1.53 1.46 1.47 1.49 1.53 1.52 1.54 1.52 3.28%
12° 1.19 1.21 1.23 1.15 1.18 1.21 1.20 1.20 1.21 4.95%
16° 8.26E-1 6.66E-1 6.54E-1 5.97E-1 6.44E-1 5.93E-1 6.28E-1 6.18E-1 6.50E-1 8.15%
18° 3.31E-1 2.92E-1 2.03E-1 2.01E-1 2.24E-1 2.22E-1 2.26E-1 2.17E-1 2.29E-1 12.23%
18.5° 2.22E-1 1.82E-1 1.50E-1 1.21E-1 1.86E-1 1.36E-1 1.53E-1 1.62E-1 1.56E-1 22.43%
19° 1.71E-1 2.08E-1 7.72E-2 7.28E-2 6.87E-2 1.99E-1 8.42E-2 9.21E-2 7.98E-2 8.77%
20° 1.70E-1 6.94E-2 1.41E-2 1.26E-2 2.66E-2 1.82E-2 1.73E-2 2.97E-2 2.30E-2 45.22%
average 6.29E-1 5.94E-1 5.41E-1 5.18E-1 5.46E-1 5.58E-1 5.47E-1 5.51E-1 5.53E-1 15.00%
@@
Figure 5: Average training and validation loss curves for MTL, MDF, RGFiL_2, MLP_2 and DAN.
@@ 4.3 Results and analysis
Figure 6: The predicted $C_{P}$ (a) and predicted errors (b) with $AoA=20$° and span width equals 28.8%. The horizontal axis denotes the x axis along the fuselage direction, and the vertical axis denotes the z axis along the fuselage normal direction.
@@
Table 5 displays the average test MSEs for MTL, MDF, RGFiL, MLP and DAN. In every row, figures in bold represent the minimum test MSEs (all the test MSEs are not normalized). Generally, the test MSEs of $C_{P}$ predicted by all models decrease as the $AoA$ increases. This phenomenon occurs because the data samples in the large $AoA$ region are densely populated, whereas the data samples in the small $AoA$ region are relatively sparse. Among RGFiL_1, RGFiL_2 and RGFiL_3, the average test MSEs of RGFiL_2 is smallest. This indicates that RGFiL show smaller prediction MSEs with $d=0.005$. Comparing RGFiL with MLP, RGFiL show smaller average test MSEs than MLP, which indicates the advantage of the multi-feature learning structure. Comparing RGFiL_2 with DAN, RGFiL_2 also exhibits smaller test MSEs than DAN. In addition, RGFiL_2 reduces the test MSEs of $C_{P}$ by an average of 15.00%, compared with DAN. The above observations indicate that RGFiL_2 is more accurate in predicting $C_{P}$ than other models.
@@
Figure 7: The predicted $C_{P}$ (a) and predicted errors (b) with $AoA=20$° and span width equals 81.8%. The horizontal and vertical axes have the same meanings as depicted in Fig. 6.
Fig. 5 depicts the average training and validation curves for all the above models in the 7-fold experiments. The fluctuations in loss variations observed in subgraphs (a), (b), (d) and (e) are larger than those in subgraph (c). This indicates a difference in the stability of the training processes (i.e., the training progress of RGFiL_2 is more stable). Furthermore, compared with subgraph (a), the models represented in subgraphs (b), (c) and (d) (MDF, RGFiL_2 and MLP_2) converge faster than MTL. These three models all incorporate geometric features as part of their inputs, which suggests that the incorporation of geometric features accelerates the convergence of their training processes. It’s worth highlighting that the models corresponding to subgraphs (c) and (d) (RGFiL_2 and MLP_2) incorporate the geometric features from additional 8 neighbors as part of their inputs, which further accelerates the convergence and further reduce the test MSEs, but at the cost of little overfitting.
Figure 8: The predicted errors with $AoA=18.5$° (a) and $AoA=7$° (b), and span width equals 15%. The horizontal and vertical axes have the same meanings as depicted in Fig. 6.
@@
Fig. 6 illustrates the predicted $C_{P}$ distributions and the corresponding prediction errors on the surface of DLR-F11 wing with $AoA=20$° and span width equals 28.8%. The DLR-F11 aircraft contains a three-stage wing, it is evident that the wing sections depicted in the graph are composed of three parts. In each row of every subgraph, the left part represents the first wing stage, the middle part represents the second wing stage, and the right part denotes the third wing stage. In subgraph (a), the $C_{P}$ distributions predicted by all models are similar, suggesting that each model is capable of learning the variations in $C_{P}$ distribution on the DLR-F11 wing surface. To clearly compare the details of the predicted $C_{P}$, we calculate
@@
where $err^{\beta}$ denotes the prediction error of the $\beta$th sample, $C_{P_{te}}^{\beta}$ denotes the real $C_{P}$ value of the $\beta$th sample in the test set, and $\hat{C}_{P}^{\beta}$ denotes the predicted value of the $\beta$th sample in the test set. The $err^{\beta}$ of each model is plotted in subgraph (b). On the first wing stage, most of the points generated by MTL and MDF show larger errors compared to the points generated by RGFiL_2. At the leading edge on the three wing stages, the points generated by MTL, MDF, MLP_2 and DNA show larger errors than the points generated by RGFiL_2. Because the leading edge of a wing is the area where its $C_{P}$ changes dramatically, RGFiL_2 has an advantage at learning $C_{P}$ variation at the leading edge of DLR-F11 wing. At the trailing edge on the lower surface of the second wing stage, the points generated by RGFiL_2 show larger errors than the points generated by MLP_2 and DAN. These observations demonstrate that RGFiL_2 is more accurate at the leading edge of wing sections than other methods.
Figure 9: The $C_{P}$ variations of test set (Real), RGFiL_2, MLP_2 and DAN with $AoA=20$°. The horizontal axis denotes the x axis along the fuselage direction, and the vertical axis denotes $C_{P}$ value.
@@
Fig. 7 illustrates the predicted $C_{P}$ distributions and the corresponding prediction errors on the surface of DLR-F11 wing with $AoA=20$° and span width equals 81.8%. Similar to Fig. 6, subgraph (a) demonstrates that all models can learn the variations of $C_{P}$ distributions. In subgraph (b), on the first wing stage, the points generated by MTL and MDF have larger errors than the points generated by RGFiL_2, MLP_2 and DAN. At the leading edge on the upper surface of the second wing stage, the points generated by MTL, MDF, MLP_2 and DAN have larger errors than the points generated by RGFiL_2. On the third wing stage, the points generated by MTL and MDF show larger errors than the points generated by RGFiL_2, MLP_2 and DAN. These observations demonstrate that RGFiL_2 is more accurate at the leading edge on the second wing stage than other methods.
@@
Figure 10: The $C_{P}$ variations of test set (Real), RGFiL_2, MLP_2 and DAN with $AoA=18.5$°. The horizontal and vertical axes have the same meanings as depicted in Fig. 9.
@@
Fig. 8 displays the prediction errors of the DLR-F11 wing section with a span width of 15% at $AoA$ of 18.5° and 7°, respectively. In subgraph (a), at the leading edge on the first wing stage, the points generated by MTL, MDF, MLP_2 and DAN show larger errors than the points generated by RGFiL_2. On the upper surface of the second stage, the points generated by MTL, MDF, MLP_2 and DAN exhibit larger errors than the points generated by RGFiL_2. On the lower surface of the second stage and both the upper surface and lower surface of the third stage, the points generated by RGFiL_2, MLP_2 and DAN are similar. In subgraph (b), it is difficult to see the differences among the models in detail. In addition, points on the upper surface of the second wing stage generated by all methods have large errors, because the DLR-F11 data samples in small $AoA$ region are sparser than those in large $AoA$. The observations in Fig. 6, Fig. 7 and Fig. 8 demonstrate that RGFiL_2 exhibits distinct advantages at the leading edge of wing sections in case with large $AoA$, while in cases with small $AoA$, RGFiL_2 closely aligns with existing methods, which is consistent with table 5.
@@
To better illustrate $C_{P}$ variations of different models, the $C_{P}$ variations of different wing sections at different $AoA$ are plotted in Fig. 9 and Fig. 10. It is worth noting that only three models (RGFiL_2, MLP_2 and DAN) are included to maintain clarity in graphs, as they exhibit relatively small prediction errors. In each subgraph, three set of curves are plotted to depict the $C_{P}$ variations on three stages of the DLR-F11 wing. The left curve represents the $C_{P}$ variation on the first wing stage, the middle curve represents the $C_{P}$ variation on the second wing stage, and the right curve represents the $C_{P}$ variation on the third wing stage. In each curve, the upper curve represents the $C_{P}$ variations on the lower surface of the DLR-F11 wing, and the lower curve represents the $C_{P}$ variations on the upper surface of the DLR-F11 wing.